# Graph Foundation Models

### v2

@@ 6 Conclusion
From the transferability principles of graphs, we review existing GFMs and ground their effectiveness from a vocabulary view to find a set of basic transferrable units across graphs and tasks. Our key perspectives can be summarized as follows: (1) Constructing a universal GFM is challenging, but domain/task-specific GFMs are approachable with the usual availability of a specific vocabulary. (2) One challenge is developing GFM following the neural scaling law, which requires principle-driven design in data engineering, architecture, and pretext tasks. This paper summarizes the current position and challenges toward the next step, which may be a blueprint for GFM to inspire relevant research.
7 Impact Statements
@@
Appendix A A collection of datasets to support large-scale pre-training
In this section, we show a collection of large-scale graph datasets from various fields to support pre-training massive-scale graph foundation models.
@@ Appendix A A collection of datasets to support large-scale pre-training
Live Graph Lab (Zhang et al., 2023d) https://livegraphlab.github.io/ A large-scale temporal graph for NFT transactions
@@ Appendix B Existing GFMs
In this section, we demonstrate existing representative GFMs and categorize them into prototype GFM, domain-specific GFM, and task-specific GFM, as shown in Table 2.
Table 2: A collection of existing GFMs
@@
Prototype GFM PRODIGY (Huang et al., 2023a)
@@
GraphText (Zhao et al., 2023b) General graph Node classification
@@
Task-specific
GFM
ULTRA (Galkin et al., 2023) Knowledge graph Knowledge graph reasoning
Appendix C A more comprehensive discussion on applying LLMs to graph
In this position paper, we mainly discuss how to build a graph-centric foundation model. Many people also understand the graph foundation model as the combination of graphs and other foundation models (Jin et al., 2023a; Li et al., 2023b; Liu et al., 2023e). In this part, we provide an introduction to those representative works. Specifically, the current integration of graph and foundational models follows two primary pathways. The first involves using graphs to augment the capabilities of other foundational models. The second employs foundational models to address challenges encountered in graph machine learning.
@@ Appendix C A more comprehensive discussion on applying LLMs to graph
The first type of work focuses on enhancing the capabilities of foundation models by graphs. Yasunaga et al. (2022a, b); Jin et al. (2023b); Xie et al. (2023) further pre-train LLMs on text-attributed graphs with a structure-aware pretext task. For example, Yasunaga et al. (2022b) trains an LLM to predict masked edges, which is formalized as pair classification on two end nodes’ attributes. Structure-aware training can effectively enhance language models’ capability on those tasks requiring structure reasoning, such as multi-hop reasoning. These works still view the graph as a second-class citizen providing auxiliary information and put more emphasis on text-centric tasks like question answering (Yasunaga et al., 2022a, b). Tang et al. (2023); Ye et al. (2023); Chai et al. (2023) improves LLMs’ capabilities tackling graph-related problems with instruction tuning. They adopt a linear map to inject the embedding generated by GNNs as a prompt token, then tune LLMs according to downstream tasks. A notable drawback of these works lies in that the tuned models don’t present capabilities beyond GNNs while introducing computation overhead.
The second line of work adopts LLMs’ capabilities to solve challenges in the graph domain. Chen et al. (2023b); He et al. (2023) adopts LLMs as encoders to augment the node attributes and enhance GNNs’ performance on node classification. Chen et al. (2023c) aims at solving the cold-start problem in node classification, and shows that actively annotating nodes with LLMs can significantly improve GNNs’ zero-shot inference capability. Zhao et al. (2023a) unifies graph and text representations of molecules via instruction-tuning, which then achieves promising zero-shot performance on property prediction.
We then provide a more comprehensive discussion on the positive potential of graphs on LLM. As discussed in Section 2.1, the straightforward application of LLms to graph learning by natural language description is less than ideal. The primary issue is that LLMs tend to approach problems with a sequentially greedy mindset (McCoy et al., 2023), which often results in superficial, ’shortcut’ solutions (Saparov & He, 2022; Dziri et al., 2023) rather than in-depth structural analysis. Thus, the principal challenge in adapting LLMs to serve as GFMs is enabling them to comprehend graph structures effectively. One promising solution involves structurally formatted prompts that LLMs can process. For instance, GraphText (Zhao et al., 2023b) introduces a tree-based prompt format that may allow LLMs to perform graph reasoning while preserving structural semantics. It remains positive potential for LLM to serve as a GFM.

### v3

@@ 6 Conclusion
From the transferability principles of graphs, we review existing GFMs and ground their effectiveness from a vocabulary view to find a set of basic transferrable units across graphs and tasks. Our key perspectives can be summarized as follows: (1) Constructing a universal GFM is challenging, but domain/task-specific GFMs are approachable with the usual availability of a specific vocabulary. (2) One challenge is developing GFMs following the neural scaling law, which requires more data collection, suitable architecture design, and properly leveraging LLMs. This paper summarizes the current position of GFMs and challenges toward the next step, which may be a blueprint for GFMs to inspire relevant research.
Acknowledgement
We want to thank Yanqiao Zhu at the University of California, Los Angeles, and Yuanqi Du at Cornell University for their constructive comments on this paper.
Haitao Mao, Zhikai Chen, Wenzhuo Tang, and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS 2246050, IIS1845081, IIS2212032, IIS2212144, IOS2107215, DUE 2234015, DRL2025244 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, the National Telecommunications and Information Administration (NTIA), the Home Depot, Amazon Faculty Award, JP Morgan Faculty Award, Microsoft Research, Meta, and SNAP. Yao Ma is supported by the National Science Foundation (NSF) under grant numbers NSF-2406648 and NSF-2406647.
Impact Statements
@@ Acknowledgement
Appendix A A Collection of Datasets to Support Pre-training
In this section, we show a collection of large-scale graph datasets from various fields to support pre-training massive-scale graph foundation models. We highly suggest using NetworkRepository (Rossi & Ahmed, 2015) for large-scale pretaining, which is the largest graph database presently available. Notably, burdensome pre-processing is required to clean those noisy and disordered data.
@@ Appendix A A Collection of Datasets to Support Pre-training
Live Graph Lab (Zhang et al., 2023e) https://livegraphlab.github.io/ A large-scale temporal graph for NFT transactions
@@
CLRS (Veličković et al., 2022) https://github.com/google-deepmind/clrs A collection of algorithmic reasoning datasets.
GraphQA (He et al., 2024) https://github.com/XiaoxinHe/G-Retriever A collection of graph question answer datasets.
@@ Appendix B Existing GFMs
In this section, we demonstrate existing representative GFMs and categorize them into primitive GFM, domain-specific GFM, and task-specific GFM, as shown in Table 2.
Table 2: A collection of existing GFMs.
@@
Primitive GFM PRODIGY (Huang et al., 2023a)
@@
LLaGA (Chen et al., 2024b)
Text-attributed graph
Node classification,
Link Prediction,
Graph classification
@@
Task-specific GFM ULTRA (Galkin et al., 2023) Knowledge graph Knowledge graph reasoning
ULTRAQUERY (Galkin et al., 2024) Knowledge graph Knowledge graph reasoning
Triplet-GMPNN (Ibarz et al., 2022) General graph algorithm reasoning
G-Retriever (He et al., 2024) General graph Graph Question Answer
GraphToken (Perozzi et al., 2024) General graph Graph Question Answer
Appendix C Practical Recipes for GFM Applications
We primarily emphasize the graph principles revolving on transferability and neural scaling law in Section 3, and 4, respectively. In this section, we provide a comprehensive application discussion with more technical details. Specifically, we introduce the feature heterogeneity issue, pretext task design, and efficiency issues in subgraph-based methods in Appendix C.1C.2, and C.3, respectively.
C.1 Tackling the Feature Heterogeneity Issue
Existing graph datasets cannot be uniformly utilized for pre-training due to the feature heterogeneity issue induced by missing features or different semantic spaces. Feature imputation techniques (Taguchi et al., 2021; Um et al., 2023; Gupta et al., 2023) are generally adapted to predict the missing attributes based on neighboring features. However, those techniques require each feature dimension to share the same semantic meaning. When features are from different semantic spaces, OFA (Liu et al., 2023b) manually converts the original features with text descriptions and then encodes the embedding with LLMs. Liu et al. (2023b) demonstrates the effectiveness and generality of using LLM embeddings to align heterogeneous node features in the text space. First, it shows that a large portion of feature heterogeneity is caused by the feature engineering process. For example, encoding text using Word2Vec (e.g. OGBN-Arxiv) and TF-IDF (e.g. Pubmed) results in different feature dimensions, leading to heterogeneity. If a unified LLM is utilized for encoding, feature heterogeneity can be solved. Second, for attributes without text attributes, OFA leverages multi-modal models to project them into textual descriptions. Specifically, OFA utilizes GIMLET (Zhao et al., 2024) to generate high-quality text descriptions for chemical molecules, and preliminarily shows that positive transferring can be achieved across diverse domains like text-attributed graphs, knowledge graphs, and molecule graphs after projecting heterogeneous features into text space. However, LLM embeddings still have limitations and their performance highly depends on the prompts provided to the LLM text encoder, remaining ample room for exploration in this area. One potential way is to borrow ideas from the CV domain. Yu et al. (2023) shows that after using LLMs to unify the feature space, further using discrete tokenization for the image can create a better latent space to further improve performance.
Feature misalignment can also be found in the inference stage between the pre-training model input and the test data. Jing et al. (2023) concatenates a learnable padding feature on the downstream task feature to align with the pre-trained GNN. However, such a technique cannot adapt to the case when the feature space is not aligned. Zhao et al. (2023a) directly abandons the original feature and utilizes the feature similarity as guidance.
C.2 Pretext Task Design
The generative self-supervised learning designs the pretext tasks (Hou et al., 2022; Hu et al., 2019; Kipf & Welling, 2016) to capture the shared data generation process among different tasks. Particularly, they attempt to predict the masked portions of the graph using the remaining structure and features. Liu et al. (2023e); Xia et al. (2023) further observe that task granularity also plays an important role in generative modeling. Specifically, employing node-level pretext tasks may lead the model to learn only low-level features (Liu et al., 2023e) while ignoring the global information essential for graph-level tasks. To address this issue, they adopt a GNN-based tokenizer to explicitly model high-level information in the pre-training stage and thus improve the downstream task performance.
More recently, the next token prediction (NTP) pretext task (Zhao et al., 2023b) achieves initial success in the molecular graph. Notably, this is the first pretext task demonstrating empirical evidence of model scaling. The potential reason for its success may be (1) the construction of a fixed token set, narrowing down the problem space in a finite set to only predict a discrete token and (2) choosing transformers as the backbone model. However, it remains unclear whether the success can be easily extended to more tasks.
C.3 Efficiency Issues in Subgraph-based Methods.
Subgraph-based extraction is a widely adopted technique in GFM to achieve inductive inference (Zeng et al., 2021) and unify different task formulations (Sun et al., 2023; Liu et al., 2023b). Nonetheless, the subgraph-based extraction leads to the following issues: (1) information loss in high-order neighborhoods; (2) duplicate sub-graph information with excessive memory consumption; (3) the time complexity of vanilla subgraph extraction grows exponentially with the number of hops, and (4) the online sub-graph sampling on the fly is also in non-acceptable inference latency (Yin et al., 2022).
Moreover, the subgraph-based method will increase the number of forward processes for a link-level task, leading to limited efficiency. Typically, for each node pair, we will extract a sub-graph based on them, and apply the forward process. Therefore, the number of forwards increases from $O(|N|)$ to $O(|E|)$, where $|N|$ and $|E|$ are the number of nodes and the number of edges. In practice, the subgraph-based method like SEAL[1] cannot be directly applied to the larger OGB-graph due to such efficiency issues. Those issues hinder the applicability of subgraph-based methods.
Designing an effective and efficient sampling method remains a major challenge in building GFM. Graph sampling techniques like (Zeng et al., 2019) and global state vectors (Fey et al., 2021) can help to alleviate these issues.
  1. 1.
    Existing GFM such as PRODIGY (Huang et al., 2023a) and OneForAll (Liu et al., 2023b) based on a subgraph-based view suffers from severe efficiency issues, especially on the link-level tasks. (1) For each node pair, those methods will extract a subgraph based on them, and apply the forward process. Therefore, the number of forwards increases from $O(|N|)$ to $O(|E|)$, where $|N|$ and $|E|$ are the number of nodes and the number of edges. (2) Moreover, the sampling subgraph may also introduce an efficiency problem (Yin et al., 2022). Subgraph-based methods sample subgraphs in either an offline or online manner. For offline sampling, they need to store subgraph patches for all possible queries, which introduces enormous memory overhead for large graphs. For online sampling, it samples subgraphs on the fly and results in non-acceptable inference latency.
  2. 2.
    To solve these efficiency issues, a potential approach is to convert GNN computing to feature precomputation (Chamberlain et al., 2022). This works for node-level and link-level tasks, but extending it to graph-level tasks is still challenging.
Appendix D Additional Principles
D.1 Principles on deeper GNN design
In section 3, we emphasize principles revolving on the transferability across datasets. Besides, another line of principles focuses on tackling the model limitation towards building effectiveness deeper GNN to capture higher-order structural information.
Principles can tell why vanilla GNNs suffer from performance degradation when increasing the number of layers and provide guidance for solutions. The principles can be majorly categorized into the following three perspectives: (i) The over-squashing problem (Topping et al., 2021) illustrates that the node representation is insensitive to information from important but distant nodes. (ii) The over-smoothing problem (Oono & Suzuki, 2019; Cai & Wang, 2020) illustrates that more aggregations lead to the node representations converging to a unique equilibrium, which loses the distinction between different nodes. (iii) The underreaching (Barceló et al., 2020) illustrates the failure to explore, cover, or affect all relevant nodes in the graph, leading to information loss. Various techniques are proposed to identify the root causes (Di Giovanni et al., 2023; Wu et al., 2023) and solve the expressiveness issues via new GNN (Yang et al., 2021) and graph transformer (Wu et al., 2022; Müller et al., 2023) architecture designs.
Despite those principles are well-studied, they can have a different position and challenges when moving from end-to-end training GNNs to the GFM requiring models to apply across different tasks and datasets. Instead of only emphasizing the effectiveness on a single dataset, building GFM raises a novel challenge for us to get good performance with a unified model backbone on diverse datasets. The GNN backbone should be able to simultaneously capture discriminative low-order neighborhood information for homophily graphs and high-order neighborhood information for heterophily graphs while the low-order ones may be noisy. Current GFMs like OneForAll (Liu et al., 2023b) empirical solve such a problem via adding virtual nodes with proper prompt designs. Nonetheless, there remains a gap between building effective and adaptive deeper GNN for GFM and the current theoretical principles.
D.2 Additional Description on the Relational Graph Vocabulary of ULTRA
[…]
Whether the relational vocabulary is suitable or not is according to the graph expressiveness theory. Typically, if two subgraphs are invariant with the same relational structure, i.e., isomorphic to the node type permutation, they will be mapping into the same token with the same representation. In contrast, if two nodes are not invariant, i.e., non-isomorphic to the node type permutation, they will be mapped into different tokens with different representations. Overall, the criterion for relational vocabulary is that two sub-structures can be mapped into the same token if and only if two sub-structures are isomorphic.
Appendix E Discussions & Open questions
E.1 More discussions on LLMs and Graphs
In this section, we provide an extended discussion on leveraging LLMs for graph-related tasks, building on the concepts introduced in Section 4.4. We provide a more comprehensive discussion on the interaction between LLMs and Graphs.
@@ E.1 More discussions on LLMs and Graphs
Specifically, the current integration of graph and foundational models follows two primary pathways. The first involves using graphs to augment the capabilities of other foundational models. The second employs foundational models to address challenges encountered in graph machine learning. The first type of work focuses on enhancing the capabilities of foundation models by graphs. Yasunaga et al. (2022a, b); Jin et al. (2023b); Xie et al. (2023) further pre-train LLMs on text-attributed graphs with a structure-aware pretext task. For example, Yasunaga et al. (2022b) trains an LLM to predict masked edges, which is formalized as pair classification on two end nodes’ attributes. Structure-aware training can effectively enhance language models’ capability on those tasks requiring structure reasoning, such as multi-hop reasoning. These works still view the graph as a second-class citizen providing auxiliary information and put more emphasis on text-centric tasks like question answering (Yasunaga et al., 2022a, b).
The second line of work adopts LLMs’ capabilities to solve challenges in the graph domain. Luo et al. (2024); Chen et al. (2024a); Wang et al. (2024b); Li et al. (2024); Ye et al. (2024) utilize the instruction fine-tuning the LLM for various capabilities including zero-shot (Li et al., 2024), link prediction (Ye et al., 2024), graph reasoning (Luo et al., 2024; Chen et al., 2024a; Wang et al., 2024b). Surprisingly, Wang et al. (2024b) observes that graph fine-tuning can even help those tasks irrelevant with graphs, e.g., mitigate hallucination, logic reasoning, and question answering. LLMs can also be utilized for graph generation (Yao et al., 2024; Wang et al., 2024c) where Wang et al. (2024c) finds that the graph generated by LLMs is biased towards more triangles and alternating 2-paths, leading to worse performance on the graph recall task.
Potential drawback on GNN-enhenced LLM Although these models can perform well, they still have two shortcomings: (1) The ability to process structures is bounded by the capabilities of GNN; (2) The instruction tuning can be costly while the tuned model can only tackle the corresponding downstream task and is not transferable to other tasks and datasets, which makes their capabilities distant from a GFM. We agree that LLM illustrates superior performance on textual node feature understanding. Nonetheless, it remains unclear whether LLM should play a key role in building GFM or just serve as a better textual feature encoder. Moreover, Stechly et al. (2023) observes that LLMs are bad at solving graph coloring instances even with multiple-round prompt. Yue et al. (2023) points out the efficiency issue of utilizing LLM on the recommendation, the downstream link prediction task. The effectiveness and efficiency of LLMs remains unclear.
E.2 Whether There exists a General Graph Vocabulary?
A shared graph vocabulary that is effective in transferability across domains and tasks remains an open question. In the current stage, we do not speculate the most ideal form of such vocabulary both across tasks and domains. Instead of transferring both across tasks and domains, the current vocabulary design can either transfer across tasks or domains. There is no unified graph vocabulary design at the current research state, as most of the graph vocabulary is either task or domain-specific, e.g., relational vocabulary in ULTRA (Galkin et al., 2023).
Despite the general graph vocabulary is challenging and not yet realized, we want to introduce one potential way toward it via graph tokenizer training, which is proposed in VQGraph (Yang et al., 2023). Specifically, it tokenizes nodes with similar structural properties into discrete codes using variants of VQ-VAE (Van Den Oord et al., 2017). After pre-training the tokenizer with a graph reconstruction objective, the discrete codes contained in the codebooks can represent typical structural patterns. The properties of the learned codes will be based on two factors: (1) the encoder and decoder architecture; and (2) the pre-training objective. The current design is still under the guidance of graph principles, remaining not generalized across all the graphs. We leave the open question whether there is a universal structure space on graph as the future work
Is it possible for GFM to transfer across different domains? For instance, can a model trained on molecular data positively transfer to KG data? The answer to this question is initially yes, where we utilize the OneForAll model (Liu et al., 2023b) as a successful showcase on cross-domain transferring. The OneForAll model unifies feature spaces from different domains by using LLM embeddings, map features, and labels into a unified text space with better transferability. Such a unified space thereby serves as the basis for GFM that can transfer across citation networks, Wikipedia knowledge graphs, and molecular graphs. In the zero-shot setting, OneForAll shows that models trained on citation networks with the node classification task can show positive transfer on molecular graphs, even surpassing the performance on foundation models specific for the science domain like Galactica (Taylor et al., 2022). We hypothesize the potential reason is the existence of transferrable patterns among domains, e.g., shared motifs. Such transferrable patterns could be modeled by the ability to recognize cycles. For example, 6-cycles are seen in molecules while 3-cycles (triangles) are critical for social networks (Granovetter, 1973). Moreover, (Ribeiro et al., 2009) indicates that there are shared patterns between the electronic circuit, the transcriptional network, and the social network despite a severe domain shift. More investigations are needed to verify whether models can effectively utilize those transferrable patterns. For the transferability between knowledge graphs and molecular graphs, we do not have empirical evidence so far. We hypothesize that if the knowledge graph involves chemistry-related knowledge, positive transferability can be achievable.
E.3 Deeper GNNs as the Backbone of GFMs
The development of deep learning generally believes the benefit from deeper Neural Network (He et al., 2016), where the worst case of deeper Neural Networks should be degraded to a shallow solution. Notably, we want to emphasize the difference between the general deeper Neural Network design and the deeper GNN design. In general deeper Neural Network design, e.g., ResNet (He et al., 2016), Transformer (Vaswani et al., 2017), a deeper Neural Network naturally leads to larger parameter scaling. However, a deeper GNN does not necessarily lead to larger parameter scaling. The key reason is that the GNN is composed of two different components including (1) the feature transformation layer and (2) the aggregation layer. Many deeper GNNs focus on increasing the non-parametric aggregation layer while the number of feature transformation layers remains small. For instance, the APPNP (Klicpera et al., 2018) on Planetoid datasets generally only utilizes two feature transformation layers with a number of parameters less than 10,000. It remains skeptical whether deeper GNNs with careful aggregation function design can achieve similar success in other domains without scaling parameter size.
E.4 Is Invariance a Necessarity for Building GFM?
We propose a graph vocabulary perspective emphasizing the invariance among graphs is essential for building Graph Foundation Model. However, it remains a mystery whether we should implicitly keep such invariance via specific Message Passing Neural Network (MPNN) design with equivariance. On the one hand, (Galkin et al., 2023, 2024) indicates the effectiveness of building GFM with equivariance. On the other hand, (Abramson et al., 2024; Wang et al., 2023c) finds that it is unnesserary to ensure invariance or equivariance with respect to global rotations and translation of the molecule. Instead, data argumentation with random rotating and translating is utilized as an implicit regularization during training. (Perozzi et al., 2024), which first encodes graph with GNNs to conduct prompts and then utilizes LLM for prediction, observes that better performance when breaking the necessary equivariance. So far there is no agree on how to preserve geometric equivariance while LLMs also demonstrate potential . In additional to geometric Neural Network design, data augmentation, loss functions, and the potential expressiveness of LLM may also provide effective solution.
E.5 Comparison with Past Relevant Literature.
Concurrent to our position paper, Jin et al. (2023a); Li et al. (2023b); Zhang et al. (2023d) reviews those methods adapting large language model (LLM) for graph, which haven’t shown transferring capabilities and thus diverge from our scope to build a graph-centric GFM. Liu et al. (2023d) further discusses existing graph pre-training and adaption techniques with a focus on their implementations. Instead of technical details, our work focuses more on the fundamental principles, e.g., geometric invariance across datasets. With principle guidance, we depict the promising and relatively elusive directions for the development of GFMs.